### v1

# thanks: Sample text

Abstract
This paper presents an overview of the Burer-Monteiro method (BM), a technique that has been applied to solve robot perception problems to certifiable optimality in real-time. BM is often used to solve semidefinite programming relaxations, which can be used to perform global optimization for non-convex perception problems. Specifically, BM leverages the low-rank structure of typical semidefinite programs to dramatically reduce the computational cost of performing optimization. This paper discusses BM in certifiable perception, with three main objectives: (i) to consolidate information from the literature into a unified presentation, (ii) to elucidate the role of the linear independence constraint qualification (LICQ), a concept not yet well-covered in certifiable perception literature, and (iii) to share practical considerations that are discussed among practitioners but not thoroughly covered in the literature. Our general aim is to offer a practical primer for applying BM towards certifiable perception.
Index Terms:
Burer-Monteiro, Semidefinite Programming, Certifiable Perception, Riemannian Staircase
I Introduction
Robotic perception is increasingly entering the world in a range of applications, from augmented reality to autonomous vehicles. These algorithms need to be both (i) real-time operable and (ii) safe and trustworthy. However, these aims are often at odds with one another, as providing rigorous guarantees on reliability often comes at substantial computational expense.
Certifiably correct algorithms – algorithms with formal guarantees on the correctness (i.e., global optimality) of their solutions – provide one means of providing reliability. However, many certifiably correct algorithms are computationally infeasible for real-time operation. This paper discusses the Burer-Monteiro method (BM), a technique that has been successfully applied to robotic perception problems to construct computationally efficient certifiably correct algorithms.
BM is a framework for solving large-scale semidefinite programming (SDP) relaxations; these SDP relaxations often arise as convex relaxations of (non-convex) perception problems. SDP relaxations are useful, as they provide strong theoretical frameworks for analyzing the optimality of solutions. However, despite being convex, standard SDP solvers do not scale well to large problems. BM provides a path to efficiently solve these SDP relaxations by exploiting the fact that solutions to these relaxations are often low-rank (as is typically the case in perception problems).
This paper is motivated by the growing amount of research in certifiably correct perception and the complex landscape of algorithms that have been developed in this area. In particular, this paper aims to convey when and why BM is suitable for certifiably correct algorithms, clarifying common points of confusion and providing a roadmap for future research.
This paper does not convey new research results, nor does it extensively cover theoretical details on BM or SDP relaxations. Instead, this paper seeks to lower the barrier to entry into this field by:
  • unifying knowledge from the literature into a single, introductory discussion of BM and certifiable perception,
  • highlighting the often unstated importance of the linear independence constraint qualification (LICQ) to a greater extent than currently found in the robotics literature, and
  • sharing practical considerations for applying BM to real-world problems.
II Notation
[…]
III The Burer-Monteiro Method (BM)
In this section we demonstrate how the Burer-Monteiro method (BM) is directly derived from standard maximum a posteriori estimation problems that arise in perception, and how BM enables efficient global optimization. Most of the material in this section is discussed across various papers in the literature (e.g., [1, 2, 3, 4]); this section concisely gathers this knowledge under a single framework. Additionally, we contain a small discussion (Section III-D) on the relevance of the linear independence constraint qualification (LICQ).
III-A From Perception to Certifiable Optimization and the BM
We begin by showing how a quadratically constrained quadratic program (QCQP) can be relaxed to a (convex) semidefinite program (SDP) via a standard technique (Shor’s relaxation [5]). We discuss the theoretical benefits conveyed by this relaxation and the computational challenges it introduces. Importantly, we discuss how the SDP relaxation provides a useful sufficient condition for the global optimality of the original QCQP.
We then introduce BM, a (non-convex) low-dimensional factorization of the SDP, to address the computational challenges of SDPs. We show how BM is naturally derived from the SDP relaxation and the benefits it provides to solving SDPs. We discuss theoretical properties of BM and how specific conditions can be evaluated to see if a locally optimal BM solution corresponds to a globally optimal solution for the corresponding SDP.
QCQPs in perception. Our paper starts by assuming that a perception problem is formulated as a quadratically constrained quadratic program (QCQP), i.e., a problem with both quadratic objective and constraint functions. Many applications in robot perception satisfy this criterion (see Section IV). In our paper, this is a non-convex QCQP (otherwise, the machinery discussed here is not necessary). For conciseness, we will consider QCQPs with only equality constraints, though the methods we discuss readily extend to inequality constraints [2]. We will consider QCQPs posed as follows:
$\min_{X\in\mathbb{R}^{n\times k}}$ $\quad\langle Q,XX^{\top}\rangle$ (1)
s.t. $\quad\langle A_{i},XX^{\top}\rangle=b_{i},\quad i=1,\ldots,m.$
$Q\in\operatorname{Sym}^{n}$ is a matrix that encodes quadratic costs, often called a data matrix in perception, and $A_{i}\in\operatorname{Sym}^{n}$ are matrices that encode quadratic constraints (which may also depend on data). The $b_{i}\in\mathbb{R}$ are constants that encode the right-hand side of the constraints. We note that the expression $\langle A,XX^{\top}\rangle$ is equivalent to the more familiar quadratic form $\operatorname{\operatorname{tr}}(X^{\top}AX)$, which simplifies to $x^{\top}Ax$ in the case of vectors.
Admittedly, much of the art of certifiable perception is in finding the right QCQP formulation for a given problem. We will not delve into this in this paper, but we point out that quadratic constraints are common in perception. E.g., the orthogonality of rotation matrices $(R^{\top}R=I)$ or the unit norm of quaternions $(q^{\top}q=1)$.
From QCQP to SDP (Shor’s relaxation). Given a QCQP of the form above, we can follow a well-known procedure to relax it to a semidefinite program (SDP). This relaxation is known as Shor’s relaxation [5]. The idea is to first introduce a variable substitution $Z=XX^{\top}$ and rewrite the QCQP as an equivalent (non-convex) SDP:
$\min_{Z\in\operatorname{Sym}^{n}}$ $\quad\langle Q,Z\rangle$ (2)
s.t. $\quad\langle A_{i},Z\rangle=b_{i},\quad i=1,\ldots,m$
$\quad Z\succeq 0,$
$\quad\operatorname{rank}(Z)\leq k$
where the implicit properties of the outer product $XX^{\top}$ are explicitly encoded as constraints on $Z$ in the form of symmetry $Z\in\operatorname{Sym}^{n}$, positive semidefiniteness $Z\succeq 0$, and rank $\operatorname{rank}(Z)\leq k$.
The sole source of non-convexity in (2) is the rank constraint [6]. By dropping the rank constraint, we obtain a convex SDP relaxation of the original QCQP.
$\min_{Z\in\operatorname{Sym}^{n}}$ $\quad\langle Q,Z\rangle$ (3)
s.t. $\quad\langle A_{i},Z\rangle=b_{i},\quad i=1,\ldots,m$
$\quad Z\succeq 0$
This relaxation is useful because it provides a sufficient condition for the global optimality of the original QCQP. Because the SDP is a relaxation of the QCQP, any solution to the SDP must lower-bound the optimal value of the QCQP. As a result, if we can find a point $X^{*}$ for the QCQP (1) such that $Z^{*}=X^{*}X^{*\top}$ solves the SDP (3), then we know that (i) the optimal value of the QCQP and the SDP are the same and (ii) $X^{*}$ is a global optimum of the QCQP. This simple idea has powerful implications. Namely, this provides a path to efficient global optimization of a non-convex problem. Rather than solving the (non-convex) QCQP directly, we can solve the (convex) SDP and try to extract a QCQP solution $X^{*}$ from the SDP solution $Z^{*}$ via e.g., singular value decomposition. However, there is a catch: the SDP relaxation is often too computationally expensive to solve with standard solvers. This computational challenge is where BM becomes useful.
The Burer-Monteiro method. BM is a technique to reduce the computational complexity of solving SDPs which have a low-rank structure, i.e., SDPs where the solution has low rank relative to the problem size: $\operatorname{rank}Z^{*}\ll n$. Fortunately, the SDP relaxations that arise from QCQPs in perception often have such low-rank structure (this can often be proven for certain noise regimes e.g., [7, 8]).
BM simply follows by introducing a factorization $Z=YY^{\top},Y\in\mathbb{R}^{n\times r}$ into the SDP relaxation. This factorization introduces two implicit constraints on $Z=YY^{\top}$: (i) positive semidefiniteness $Z\succeq 0$ and (ii) low-rank $\operatorname{rank}(Z)\leq r$. The resulting BM problem is,
$\min_{Y\in\mathbb{R}^{n\times r}}$ $\quad\langle Q,YY^{\top}\rangle$ (4)
s.t. $\quad\langle A_{i},YY^{\top}\rangle=b_{i},\quad i=1,\ldots,m$
With $r\ll n$, the BM problem has a much lower-dimensional state space than the original SDP. This reduces the number of computational operations required to solve the problem, making it more tractable. Modern BM solvers can operate on problems with $n$ in the tens of thousands on standard laptops in seconds. This is in stark contrast to generic SDP solvers, which on similar machines can struggle with problems of size $n$ in the hundreds [9].
Formulation (4) may look familiar; this is effectively the same factorization we used to relax the QCQP to the SDP in the first place. Specifically, when the BM variable $Y$ has the same number of columns as $X$ in the original QCQP (i.e., $r=k$), the BM problem is equivalent to the original QCQP. In general, the BM problem is also a non-convex QCQP. This may seem like a circular way to arrive at the original problem, but this circuitous route provides us two key viewpoints.
First, we can see that by increasing the rank $r$ of the BM problem, we can view BM as providing a hierarchy of relaxations of the original problem. Intuitively, increasing $r$ increases the free dimensions that an optimizer may access, allowing for new descent directions that can avoid what are local minima in lower-rank relaxations of the BM problem.
The second advantage is the capacity for efficient global optimization. This stems from connection to the SDP relaxation. Formally, the following relationship holds:
$f^{*}_{\text{SDP}}\leq f^{*}_{\text{BM}}\leq f^{*}_{\text{QCQP}}.$ (5)
Where $f^{*}_{\text{SDP}}$, $f^{*}_{\text{BM}}$, and $f^{*}_{\text{QCQP}}$ are the optimal values of the SDP, BM, and QCQP problems, respectively and the BM problem is not lower-dimensional than the original QCQP (i.e., $r\geq k$). This relationship provides a sufficient condition to certify the global optimality of BM and QCQP solutions in the case when the inequalities are tight (i.e., $f^{*}_{\text{SDP}}=f^{*}_{\text{BM}}=f^{*}_{\text{QCQP}}$).
With this in mind, we can solve a series of BM problems with increasing rank $r$ until a solution $Y^{*}$ is found that is also a low-rank factorization of the SDP solution $Z^{*}=Y^{*}Y^{*\top}$, an approach known as the Riemannian Staircase [10] when the BM problems are solved using Riemannian optimization. As we have a solution to the SDP, we have $f_{\text{SDP}}$. Because $Y^{*}$ attains the same objective value as $Z^{*}$, from the relationship in (5), we also know we have a globally optimal solution to the BM problem with objective value $f_{\text{BM}}=f_{\text{SDP}}$. Finally, if $Y^{*}$ has the same rank as the dimension of the QCQP (1) (i.e., $\operatorname{rank}(Y^{*})=k$), then we have a globally optimal solution to the original QCQP with objective value $f_{\text{QCQP}}=f_{\text{BM}}=f_{\text{SDP}}$.111 We can always obtain a certifiably optimal BM solution $Y^{*}$. However, it is possible that the SDP relaxation is not tight (i.e., $f^{*}_{\text{SDP}}<f^{*}_{\text{QCQP}}$ and $\operatorname{rank}(Y^{*})>k$); in this case, QCQP solutions cannot be certified but (5) can be used to bound the QCQP solution’s suboptimality. Additionally, an approximate solution to the QCQP can be extracted from the BM solution $Y^{*}$ via singular value decomposition, subsequently projected to the QCQP’s feasible set, and used as a starting point for further optimization. This has been used to great effect in practice (e.g., [4]).
To take advantage of this theoretical framework, we need two algorithmic tools: (i) a way to perform optimization on BM problems and (ii) a way to certify whether a BM solution maps to an SDP solution. As we are concerned with runtime, each of these items must be computationally efficient. We discuss these two items in the following sections.
III-B Local Solvers Compatible with BM
In this subsection, we review optimization algorithms that are suitable for identifying local solutions to the intermediate problems (4) introduced by the BM hierarchy. These solvers are not limited to BM – they can be applied to a wide range of optimization problems beyond the specific structure of (4). However, the choice of solver has a major impact on the efficiency and reliability of the overall BM approach. We highlight the most relevant solvers below.
One approach is to view and (locally) solve (4) as an instance of a generic nonlinear program [1]. We refer to this as an extrinsic approach, because it enforces the search space $\mathcal{M}\triangleq\{Y\in\mathbb{R}^{n\times r}:\langle A_{i},YY^{\top}\rangle=b_{i},i=1,\ldots,m\}$ using explicit constraints in the ambient Euclidean space $\mathcal{E}\triangleq\mathbb{R}^{n\times r}$. The original work of Burer and Monteiro [1] uses an extrinsic method to locally solve (4), which enforces the search space constraints by optimizing over the augmented Lagrangian function. Recent work [2] extends BM to generic low-rank SDPs with inequality constraints and also adopts an extrinsic local solver implemented in IPOPT [11].
Intrinsic (Riemannian) solvers. In contrast to the extrinsic local solvers above, more recent works further exploit the geometric structure of (4) through an intrinsic perspective. In practice, for many robot perception applications, $\mathcal{M}$ turns out to be “standard” matrix manifolds whose geometries are well studied (e.g., a Stiefel manifold [7, 12, 13]). As such, existing theories and implementations of Riemannian optimization [14, 15] directly apply. These approaches solve (4) by operating on the manifold intrinsically. The intrinsic approach is favorable because the corresponding Riemannian optimization problem is unconstrained. This enables the use of unconstrained optimization algorithms (generalized to operate on manifolds) that by design produce a sequence of feasible iterates, enjoy convergence guarantees similar to those of extrinsic solvers, and have empirically shown to be substantially more efficient than extrinsic solvers in perception applications, e.g., for pose graph optimization [2].
For typical instances of (4) arising from robot perception applications, second-order (Newton-type) solvers combined with globalization strategies (e.g., trust-region) have proven particularly effective. The second-order property of the solver helps to evade spurious first-order critical points (where the gradient vanishes) and achieves a fast (superlinear) local convergence rate. The globalization strategy further prevents possible divergence (which is possible with the vanilla Newton’s method), and ensures optimization converges from any initial guess. A prominent example that follows this design principle is the Riemannian trust-region (RTR) algorithm [16] used by many state-of-the-art certifiable methods [7, 12, 4]. At every iteration, RTR approximately minimizes a local second-order model of (4) under a trust-region constraint, which limits the magnitude of the computed update. The size of the trust region is adjusted dynamically, so that it acts as a safeguard when the quality of the model function is poor, but still does not interfere with the fast local convergence of typical second-order optimization. The model minimization is typically done iteratively using the truncated conjugate gradient (tCG) method (e.g., see [15, Section 6.3] for details). Closely related to RTR is the Riemannian Levenberg-Marquardt (LM) method [14, Section 8.4.2]. LM uses the same local quadratic model of the objective as the Gauss-Newton method. Instead of changing the trust-region size as in RTR, LM dynamically adjusts a regularization term that is added to the model function, which can be interpreted as the Lagrangian form of a trust-region constraint and plays a similar role of discouraging large updates. Among recent certifiable perception methods, the rotation averaging method by Dellaert et al. [13] uses LM as implemented in GTSAM [17] to solve BM problems defined on rotation groups with increasing dimensions.
Extensions to distributed/parallel computing. To extend the Riemannian Staircase approach [10, 7] to the distributed regime, Tian et al. [3] developed Riemannian Block Coordinate Descent (RBCD) as a distributed local optimization method that leverages the product manifold structure that naturally arises in many robot perception tasks. Additional enhancements to this idea, including an extension to operate under asynchronous communication, are proposed in [18, 19]. In general, many other distributed optimization algorithms are theoretically compatible with BM if they can identify first-order critical points of (4). Examples include methods based on distributed Riemannian gradient descent [20, 21, 22]. Recent works have also extended the alternating direction method of multipliers (ADMM) to solve distributed optimization over factor graphs [23, 24, 25], although formal convergence guarantees remain to be explored under Riemannian manifold constraints. Fan and Murphey [26] developed an extrinsic method based on accelerated majorization minimization for distributed pose graph optimization.
III-C Certification
Burer-Monteiro SDP
Stationarity $S_{\lambda}Y=\bf{0}$ $S_{\lambda}Z=\bf{0}$
Primal feasibility $\langle A_{i},YY^{\top}\rangle=b_{i}$ $\langle A_{i},Z\rangle=b_{i}$
$Z\succeq 0$
Dual feasibility not relevant $S_{\lambda}\succeq 0$
TABLE I: Optimality conditions for the BM and SDP problems, where $S_{\lambda}\triangleq Q+\sum\lambda_{i}A_{i}$. The conditions for BM are simply necessary for first-order optimality, while the conditions for the SDP are necessary and sufficient for global optimality. Dual feasibility in BM is not relevant, as it is not needed to construct the certification scheme we describe. Both sets of conditions can be found from the KKT conditions of the respective problems [2].
As previously mentioned, a key capability is certifying whether a BM solution $Y^{*}$ is a low-rank factor for a solution $Z^{*}=Y^{*}(Y^{*})^{\top}$ of the original SDP. If so, the BM solution is guaranteed to be globally optimal for the BM problem $f_{\text{BM}}=f_{\text{SDP}}$), and we now have a certified lower bound on the attainable cost of the original QCQP ($f_{\text{BM}}\leq f_{\text{QCQP}}$). This lower bound allows for certification of a QCQP solution when the SDP relaxation is tight ($f_{\text{QCQP}}=f_{\text{BM}}=f_{\text{SDP}}$), which is often the case in practice.
The naive approach to certifying a BM solution is to generate the corresponding SDP solution $Z^{*}=Y^{*}(Y^{*})^{\top}$ and check whether $Z^{*}$ satisfies the Karush-Kuhn-Tucker (KKT) conditions of the SDP relaxation, which are necessary and sufficient for optimality of the SDP relaxation [27]. However, $Z^{*}$ would be generically dense, and thus incur substantial computational overhead. We instead describe a separate approach (described in [2]) that leverages the low-dimensional BM factorization $Y^{*}$.
As described in Table I, comparison of the KKT conditions of the BM and SDP problems reveals that a first-order stationary point of the BM problem $Y^{*}$ can generate a candidate solution to the SDP problem $Z^{*}=Y^{*}(Y^{*})^{\top}$, which automatically satisfies all SDP optimality conditions except for the dual feasibility condition $S_{\lambda}\succeq 0$. As a result, a BM solution can be certified as globally optimal by checking positive semidefiniteness of the certificate matrix,
$S_{\lambda}=Q+\sum_{i=1}^{m}\lambda_{i}A_{i},$ (6)
where $Q$ is the data matrix describing the original QCQP, and $\lambda_{i}$ and $A_{i}$ are the Lagrange multipliers and constraint matrices. In practice, the fastest and most reliable way to evaluate $S_{\lambda}\succeq 0$ is to compute a Cholesky factorization of $S_{\lambda}+\epsilon I$ for a small $\epsilon>0$, which will fail if $S_{\lambda}$ is not positive semidefinite.222 If $S_{\lambda}$ is not positive semidefinite, then a negative eigenpair of $S_{\lambda}$ can be used to construct a second-order descent direction to kickstart optimization at the next level ($r+1$) of the BM hierarchy. See e.g., [10] for more details.
To actually compute the Lagrange multipliers $\lambda$ at a candidate solution $Y^{*}$, one can use the BM stationarity condition from Table I which is equivalent to,
$\sum_{i=1}^{m}(A_{i}Y^{*})\lambda_{i}=-(QY^{*}),$ (7)
and solve a linear system for $\lambda$. Given $\lambda$, the certificate matrix $S_{\lambda}$ can be computed and $S_{\lambda}\succeq 0$ can be checked.
III-D The Role of the LICQ in BM
The linear independence constraint qualification (LICQ) is a standard constraint qualification that plays a pivotal role in the success of the BM framework and is prevalent in many robot perception applications (Section IV). This subsection briefly discusses this topic for interested practitioners.
The LICQ is satisfied if the gradients of the constraints are linearly independent. In the context of BM, this means that $\{\nabla\langle A_{i},XX^{\top}\rangle\mid i=1,\ldots,m\}$ is a linearly independent set.
Local optimization (Section III-B). There are two important aspects of the LICQ with respect to local optimization: (i) the LICQ is closely connected to the use of Riemannian optimization and (ii) the LICQ is tightly connected to the convergence of many local optimization algorithms.
Regarding Riemannian optimization, if the LICQ is satisfied globally (i.e., at all feasible points) for the BM problem (4), then the search space $\mathcal{M}$ of (4) forms a smooth manifold [15, Ch. 7]. However, it is important to note that the LICQ alone is not enough for the practical success of Riemannian optimization, and additional information on the knowledge of $\mathcal{M}$ is needed to have efficient numerical implementations (more specifically, to implement the retraction operators [28, Ch. 3] within Riemannian optimization).
Regarding the convergence of local solvers, the LICQ is a key ingredient in establishing efficient convergence for many general-purpose optimization algorithms (e.g., interior-point methods [29, Ch. 19.8]). The LICQ (along with the second order sufficiency condition) is necessary to guarantee nonsingularity of the primal-dual KKT system matrix, which is key in establishing superlinear convergence of second-order and Newton-type methods. This dependence on the LICQ highlights the challenges of efficiently solving the BM problem in its extrinsic form when the LICQ is not satisfied. However, this does not preclude the use of methods which do not rely on the LICQ (e.g., penalty methods), but may not be as efficient as those that do rely on the LICQ.
Certification (Section III-C). The LICQ is key to performing efficient certification. This is because it is the weakest condition that is necessary and sufficient for the existence of unique Lagrange multipliers [30]. Recall that the certificate matrix $S_{\lambda}$ in (6) depends on the Lagrange multipliers $\lambda$, which, at a candidate solution $Y^{*}$, are determined by the stationarity condition (7). However, this linear system only admits a unique solution $\lambda^{*}$ if the LICQ holds [30].
If the LICQ is satisfied, certification is done by solving the linear system (7) for $\lambda^{*}$, then forming $S_{\lambda}$, and finally evaluating positive semidefiniteness $S_{\lambda}\succeq 0$. If the LICQ is not satisfied, then it is possible for there to exist many different $\lambda^{*}$ that satisfy (7), but only some of which may correspond to a positive semidefinite certificate matrix $S_{\lambda}\succeq 0$. Without the LICQ, certification is equivalent to finding an intersection of the affine space defined by (7) and the positive semidefinite cone, a semidefinite feasibility problem in its own right and generally as expensive to solve as the original SDP (3).
IV Applications
To date, BM has been applied to a variety of perception tasks, including rotation and pose synchronization [7, 12, 31, 13], landmark-based SLAM [32], range-aided SLAM [4], sensor network localization [33], essential matrix estimation for structure from motion [34], and semantic segmentation via Markov random fields [35]. In this section we discuss commonalities across existing BM applications to try to understand the where and why of BM’s success in perception.
Of interest is that all of these applications leveraged well-studied Riemannian manifolds (e.g., the Stiefel manifold or the unit sphere) in formulating their problems and solved them in their intrinsic forms via Riemannian optimization. The only instance in the perception literature that we are aware of which used a purely extrinsic solver for BM is [2], which was done to compare the performance of extrinsic and intrinsic solvers in the context of pose-graph optimization. Additionally, Karimian and Tron [34] posed essential matrix estimation as optimization over the Stiefel manifold with an additional constraint (to represent epipolar geometry), though still used intrinsic descriptions of the problem.
Considering this preference, it is natural to ask why have no perception problems used an extrinsic formulation for local optimization? This is interesting, as arriving at an extrinsic BM formulation requires less work; intrinsic formulations require identifying manifold structure within the BM extrinsic formulation (4) and defining additional manifold notions (e.g., a retraction operator). There are several explanations for this preference towards intrinsic formulations despite the additional work required. We posit that this preference is due to a combination of: common problem structure, bias towards successful formulations, and the availability of optimization software.
Geometric structure. Perception problems typically possess smooth, geometric structure that is naturally expressed as well-studied manifolds. For example, orthogonality and unit-norm constraints appear throughout perception – correspondingly the Stiefel manifold is ubiquitous in certifiable perception. In fact, all existing works using BM for certifiable perception can be expressed as optimization over the Stiefel manifold and Euclidean space.333 Note that the unit-sphere is a special case of the Stiefel manifold. Karimian and Tron [34] are a notable partial exception, as they derive a custom manifold by adding an additional explicit constraint to the Stiefel manifold.
Bias towards successful formulations. There are often many different QCQP formulations of the same problem, which typically differ in the tightness of their SDP relaxations. While understanding the relationship between a QCQP and the tightness of its SDP relaxation is an open area of research (e.g., [8]), we have empirically found that (a) many possible formulations of perception problems are not tight, and (b) the Stiefel manifold often leads to tight SDP relaxations. As previously noted, all works to date have used some formulation that can be related to (special cases of) the Stiefel manifold. Therefore, this apparent bias towards intrinsic formulations may be “natural selection” appearing due to ideal properties of the Stiefel manifold.
Available optimizers. Existing manifold optimization libraries are relatively mature, allowing for straightforward evaluation of intrinsic formulations without requiring the user to implement their own optimizer. In contrast, the apparent lack of a “standard” extrinsic solver for BM represents a barrier to evaluating more general extrinsic formulations in practice. We review available solvers in Section IV.
Additionally, as noted in Section III-B, Riemannian optimization conveys substantial computational benefits over extrinsic optimization. The combined advantages in reliability and efficiency of Riemannian solvers further incentivize the additional effort required to formulate problems intrinsically.
There are several practically oriented considerations in applying BM to certifiable perception problems. We base these considerations on our own experiences as practitioners. We focus on numerical conditioning, sparsity, and existing solvers. The first two points (conditioning and sparsity) are particularly relevant for problems such as SLAM, which often manifest as large-scale estimation over sparse graphs. The final point (existing solvers) is useful for all practitioners, as there are many flexible and performant optimization libraries available.
Numerical (pre)conditioning for RTR. Many perception problems lead to large and ill-conditioned optimization problems, which presents a substantial challenge for many optimization algorithms. In the following, we focus on RTR, a widely used local solver for BM (Section III-B) whose performance is highly dependent on the conditioning of the trust-region subproblem that is solved at each iteration.
Intuitively, preconditioning in the case of RTR attempts to transform the trust-region subproblem’s loss-landscape from highly elongated to spherical (i.e., isotropic), allowing for more efficient iterations towards the solution. In practical implementations, preconditioning is often carried out by transforming the current search direction (e.g., provided by the negative gradient) via a symmetric and positive definite map $P$. For perception applications, which are often poorly conditioned, we have observed that a suitable preconditioner is often indispensable for RTR to obtain an acceptably accurate solution within the runtime constraints of real-time robotics. In particular, choosing a good preconditioner has often led to several orders of magnitude improvements in terms of (i) number of tCG iterations to converge to a suitable trust-region subproblem solution, and (ii) overall runtime.
There is a rich literature and theory behind the construction and analysis of preconditioners (e.g., [36, Ch. 10] and [37, Ch. 10.2.7]), which we do not delve into here. We instead outline relevant considerations and suggest a generally successful preconditioner for the RTR algorithm applied to BM problems.
Important considerations: There are roughly three aspects in which a preconditioner affects the runtime of an optimization algorithm: (i) the cost of calculating the preconditioner, (ii) the cost of applying the preconditioner, and (iii) the savings in the number of iterations required to converge. Ideally a preconditioner is computed once and reused across many iterations, amortizing the cost of computation. In general, there is no one-size-fits-all preconditioner, and the choice of preconditioner depends on the problem structure.
Preconditioner for RTR: For RTR, an ideal preconditioner approximates the inverse of the Riemannian Hessian of the cost function. The Riemannian Hessian depends on the point on the manifold at which it is evaluated, and thus typically changes at each iteration. Importantly, the Riemannian Hessian is closely related to the Euclidean Hessian [15, Ch. 5]. Furthermore, for the BM formulation we presented (4), the Euclidean Hessian is exactly the data matrix $Q$ and is therefore constant. As a result, $Q^{-1}$ appears as a natural preconditioner candidate, for it is closely related to the Riemannian Hessian and can be computed once and used repeatedly. However, preconditioners must be positive definite. Fortunately, in the problems we have encountered, $Q$ has been positive semidefinite and thus becomes positive definite with a small regularization $Q+\mu I$.
As this would suggest, in the problems we have seen, the inverse of the regularized data matrix $P=(Q+\mu I)^{-1}$ has been a successful preconditioner for the RTR trust-region subproblem. The regularization term $\mu\in\mathbb{R}$ is typically chosen to keep $P$’s condition number below $10^{6}$. Instead of directly computing the inverse, a Cholesky factorization $R^{\top}R=Q+\mu I$ is computed (with a sparsity-promoting ordering). This allows the preconditioner to be applied with greater numerical stability and efficiency via forward- and back-substitution with the Cholesky factor $R$.
Sparsity. The data and constraint matrices of many problems are often sparse (many elements are zero). For large-scale problems, exploiting this sparsity can lead to significant computational savings. Oftentimes, exploiting sparsity involves using specific data structures, software libraries, and algorithms that are designed to handle sparse matrices efficiently. Sparsity reduces memory footprint, improves computational efficiency, and can lead to more numerically stable procedures. In distributed settings, sparsity promotes communication efficiency. Most state-of-the-art distributed optimization methods (e.g., those in Section III-B) seek to preserve and leverage sparsity, so that robots only exchange information over a small number of variables that couple together robots’ local factor graphs.
Existing solvers. As previously mentioned, there are a number of general-purpose Riemannian optimization libraries that have been used for certifiable perception. The Manopt family [28, 38, 39] spans MATLAB, Python, and Julia. In C++, there is ROPTLIB [40], GTSAM [17], and the Optimization library by Rosen [41]. Additionally, while, to our knowledge, Ceres [42] and g2o [43] have not been used for certifiable perception, they could also be used to solve BM problems in the intrinsic form. In the extrinsic setting, to the best of our knowledge, no standard BM solvers exist, although practitioners have written custom interfaces to more general optimization libraries (e.g., Rosen [2] used IPOPT [11]).
V Open directions
BM, and certifiable perception more broadly, has shown great promise in advancing robotic capabilities. We foresee several exciting frontiers for future research along these lines. While we focus on BM, aspects of these directions are also broadly relevant to semidefinite optimization, certifiable perception, and general optimization.
Tools to improve accessibility. While BM has been applied to a variety of problems, each new problem requires a bespoke formulation. Obtaining these formulations typically necessitates substantial algebraic manipulation and an in-depth understanding of the underlying theory behind BM and certifiable perception. This represents a significant barrier to entry for practitioners. It is unknown whether methodologies could be developed to automatically derive useful QCQP formulations (i.e., that satisfy the LICQ and possess tight SDP relaxations). Such a tool could possibly leverage the growing catalogue of successful formulations or develop an approach for finding QCQP formulations that approximate a given problem. Additionally, it may be possible that no such useful QCQP formulations could be found for a given problem; an impactful tool in this case could assist a user in determining if such a formulation is likely to be found. Advances in this direction would greatly benefit from deep understanding of structural relationships between QCQPs and their SDP relaxations.
Robust costs and outlier rejection. Outlier rejection (via robust cost functions) has been explored in certifiable perception [44], however every formulation to date has had to introduce redundant constraints and, as a result, violate the LICQ. Without the LICQ these formulations have not been able to leverage the computational benefits of BM. It is an open question whether it is possible to formulate robust cost functions that preserve the LICQ.
Distributed optimization. In centralized settings, practitioners have largely converged to a set of well-implemented trust-region algorithms (such as RTR and Levenberg-Marquardt). In the distributed setup, despite the initial progress discussed in earlier sections of this paper, there is still no such consensus and additional research is required especially when considering limited communication (either due to bandwidth restrictions or privacy concerns). We believe this opens the possibility of contributions in: (i) designing distributed algorithms for optimization and certification, (ii) analyzing the convergence of these algorithms, and (iii) developing standard software tools.
VI Conclusion
We have presented an introductory overview of the Burer-Monteiro method (BM) for certifiable perception problems. We discussed key theoretical properties, outlined important theoretical requirements that are not typically discussed in the literature, and provided practical considerations for applying BM to perception problems. We also discussed open directions for future research in this area.
We believe that BM will play an important role in the future of robotics and perception. However, we also believe that BM is not a one-size-fits-all solution. BM should be used when both of the following conditions are met: (i) tight SDP relaxations can be constructed via Shor’s relaxation and (ii) the QCQP formulation globally satisfies the LICQ. However, many perception problems will likely not be able to satisfy these conditions. Indeed, better understanding the boundary of BM’s applicability and determining when alternative approaches (e.g., [45, 46]) are important questions that demand more practical insights and theoretical investigations.
VII Acknowledgements
This work was supported in part by ARL DCIST under Cooperative Agreement Number W911NF-17-20181, by ONR under BRC Award N000141712072, by ONR grants N00014-23-12164, and N00014-19-1-2571 (Neuroautonomy MURI), the MIT Portugal Program, and the MIT Lincoln Laboratory Autonomous Systems Line which is funded by the Under Secretary of Defense for Research and Engineering through Air Force Contract No. FA8702-15-D-0001, and MIT Lincoln Laboratory’s Certifiable Semialgebraic Optimization program under contract FA8702-15-D-0001.